# How many measurements are enough? Bayesian recovery in inverse problems with general distributions

### v1

@@ C.2.2 Disjointly-supported measures induce well-separated measurement distributions
The following lemma pertains to the pushforwards of measures supported in $\mathbb{R}^{n}$ via the forwards operator $A$ and noise $e$. Specifically, it states that if two distributions $\mathcal{P}_{\mathsf{int}}$ and $\mathcal{P}_{\mathsf{ext}}$ are disjointly supported then their corresponding pushforwards $\mathcal{H}_{\mathsf{int},A}$ and $\mathcal{H}_{\mathsf{ext},A}$ are, on average with respect to $A\sim\mathcal{A}$, well-separated, in the sense of their TV-distance. It is generalization of [47, Lem. 3.2] that allows for arbitrary distributions $\mathcal{A}$ of the forward operators, as opposed to just distributions of Gaussian random matrices.
@@
$\mathbb{E}_{x\sim\mathcal{P}_{\mathsf{ext}}}[\mathbb{P}_{A\sim\mathcal{A}}({\|Ax-A\tilde{x}\|}<2\sqrt{c}(\eta+\sigma)]$ $\leq\mathbb{E}_{x\sim\mathcal{P}_{\mathsf{ext}}}\left[\mathbb{P}_{A\sim\mathcal{A}}\left({\|Ax-A\tilde{x}\|}<\frac{2}{\sqrt{c}}{\|x-\tilde{x}\|}\right)\right].$
@@
$\mathbb{E}_{x\sim\mathcal{P}_{\mathsf{int}}}[\mathbb{P}_{A\sim\mathcal{A}}({\|A(x-\tilde{x})\|}\geq\frac{\sqrt{c}}{2}(\eta+\sigma))]\leq\mathbb{E}_{x\sim\mathcal{P}_{\mathsf{int}}}\left[\mathbb{P}_{A\sim\mathcal{A}}({\|A(x-\tilde{x})\|}\geq\frac{\sqrt{c}}{2}{\|x-\tilde{x}\|})\right]$
@@ C.2.3 Replacing the real distribution with the approximate distribution
We next establish a result that allows one to upper bound the failure probability based on draws from the real distribution $\mathcal{R}$ with the failure probability based on draws from the approximate distribution $\mathcal{P}$. This lemma is a key technical step that aligns the prior distribution with the posterior. The specific bound is given in terms of the Wasserstein distance between $\mathcal{R}$ and $\mathcal{P}$ and several of the concentration bounds defined in §2. This is a significant generalization of [47, Lem. 3.3] that allows for arbitrary distributions $\mathcal{A}$, $\mathcal{E}$ for the forwards operator and noise.
@@
Let $\varepsilon,\sigma,d,t\geq 0$, $c\geq 1$, $\mathcal{E}$ be a distribution on $\mathbb{R}^{m}$ and $\mathcal{R},\mathcal{P}$ be distributions on $\mathbb{R}^{n}$ such that $W_{\infty}(\mathcal{R},\mathcal{P})\leq\varepsilon$. Let $\Pi$ ben an $W_{\infty}$-optimal coupling of $\mathcal{R}$ and $\mathcal{P}$ and define the set $D=\{x^{*}-z^{*}:(x^{*},z^{*})\in\mathrm{supp}(\Pi)\}$. Let
@@
$q=\mathbb{P}_{z^{*}\sim\mathcal{R},A\sim\mathcal{A},e\sim\mathcal{E},\hat{z}\sim\mathcal{P}(\cdot|Az^{*}+e,A)}[{\|z^{*}-\hat{x}\|}\geq d].$
@@
$\begin{split}p&=\mathbb{P}_{x^{*}\sim\mathcal{R},A\sim\mathcal{A},e\sim\mathcal{E},\hat{x}\sim\mathcal{P}(\cdot|Ax^{*}+e,A)}[x^{*}\in B_{1,\hat{x}}]\\ q&=\mathbb{P}_{z^{*}\sim\mathcal{P},A\sim\mathcal{A},e\sim\mathcal{E},\hat{z}\sim\mathcal{P}(\cdot|Az^{*}+e,A)}\left[z^{*}\in B_{2,\hat{z}}\right]\end{split}.$ (C.10)
@@
×$\times\int 1_{B_{2,\hat{x}}}(z^{*})\,\mathrm{d}\mathcal{P}(\cdot|Az^{*}+e^{\prime},A)(\hat{x})\,\mathrm{d}e^{\prime}\,\mathrm{d}\mathcal{A}(A)\,\mathrm{d}\Pi(x^{*},z^{*}).$
@@ C.2.4 Decomposing distributions
The following lemma is in large part similar to [47, Lem.  A.1]. However, we streamline and rewrite its proof for clarity and completeness, fix a number of small issues and make an addition to the statement (concerning the set $S$ the distribution in whose support it is contained) which is important for proving our main result.
@@
then there exist distributions ,′′$\mathcal{R}^{\prime},\mathcal{R}^{\prime\prime},\mathcal{P}^{\prime},\mathcal{P}^{\prime\prime}$, a constant $0<\delta^{\prime}\leq\delta$ and a discrete distribution $\mathcal{Q}$ with $\mathrm{supp}(\mathcal{Q})=S$ satisfying
@@
This lemma states that two distributions that are close in Wasserstein $p$-distance, and for which at least one has small approximate covering number (C.15), can be decomposed into mixtures (iii) of distributions, where the following holds. One of the distributions, say $\mathcal{P}^{\prime}$, is close (i) in Wasserstein-$\infty$ distance to a discrete distribution $\mathcal{Q}$ with the cardinality of its support (iv) bounded by the approximate covering number. The other $\mathcal{R}^{\prime}$ is close in Wasserstein-$\infty$ distance to $\mathcal{P}^{\prime}$. Moreover, both mixtures (iii) are dominated by these distributions: the ‘remainder’ terms $\mathcal{P}^{\prime\prime}$ and $\mathcal{R}^{\prime\prime}$ associated with a small constant $\delta^{\prime}\leq\delta$, meaning they are sampled with probability $\leq\delta$ when drawing from either $\mathcal{P}$ or $\mathcal{R}$. Note that if $p<\infty$ then the Wasserstein-$\infty$ distance between $\mathcal{R}^{\prime}$ and $\mathcal{P}^{\prime}$ may get larger as $\delta$ shrinks, i.e., as the remainder gets smaller. However, this does not occur when $p=\infty$, as (ii) is independent of $\delta$ in this case.
@@
1. Construction of $\mathcal{Q}^{\prime}$. We will now define a finite measure $\mathcal{Q}^{\prime}$. The point of $\mathcal{Q}^{\prime}$ is to, concentrate the mass of the measure $\mathcal{P}$ into the centres of the balls $u_{i}$. If the sets $B(u_{i},\eta)$ are disjoint, then this is straightforward. However, to ensure that $\mathcal{Q}^{\prime}$ is indeed a probability measure, we need to normalize and account for any non-trivial intersections. This is done via the function $f$. Pick some arbitrary $\hat{u}\notin\{u_{1},\ldots,u_{l}\}$ and define
@@
$\Pi(E,\mathbb{R}^{n})=\mathcal{P}\left(E\cap\bigcup^{l}_{i=1}B(u_{i},\eta)\right)+\mathcal{P}\left(E\backslash\bigcup^{l}_{i=1}B(u_{i},\eta)\right)\equiv\mathcal{P}(E),$
@@
$\sum_{i=1}^{l}\Pi(U_{i},u_{i})\mathcal{Q}^{\prime}(u_{i})+\Pi(\hat{U},\hat{u})\mathcal{Q}^{\prime}(\hat{u})=\Pi(\hat{U},\hat{u})\mathcal{Q}^{\prime}(\hat{u})$
@@
We have measures $\mathcal{P}$, $\mathcal{Q}^{\prime}$ and $\mathcal{R}$ and couplings $\Pi$ of $\mathcal{P},\mathcal{Q}^{\prime}$ and $\Gamma$ of $\mathcal{P}$, $\mathcal{R}$. We will in a sense, couple $\Pi,\Gamma$. Since $(\mathbb{R}^{n})^{3}$ is a Polish space, by [9, Lem. 8.4], there exists a coupling $\Omega$ with
@@
$E_{1}$ :=$:=\{(x_{1},x_{2},x_{3}):{\|x_{1}-x_{2}\|}>\eta,z\in\mathbb{R}^{n}\},$
$E_{2}$ :=$:=\{(x_{1},x_{2},x_{3}):{\|x_{1}-x_{3}\|}>\rho/\delta^{1/p},y\in\mathbb{R}^{n}\},$
@@
$\mathcal{R}^{\prime}(A)$ $=\Omega(\mathbb{R}^{n},\mathbb{R}^{n},A|E^{c}),$
@@
We now prove Theorem 3.1. This follows a similar approach to that of [47, Thm. 3.4], but with a series of significant modifications to account for the substantially more general setup considered in this work. We also streamline the proof and clarify a number of key steps.
@@
Then it is readily checked that $\Gamma_{\tilde{z}}(\cdot)$ defines a probability measure. Note also that $\Gamma_{\tilde{z}}$ is supported on a ball of radius $\eta^{\prime}$ around $\tilde{z}$, since $\mathrm{esssup}_{\Gamma}({\|x-y\|})\leq\eta^{\prime}$. Recall that $\Gamma$ is a coupling between $\mathcal{P}^{\prime}$ and $\mathcal{Q}$. Let $E\subseteq\mathbb{R}^{n}$ be a Borel set. Then Lemma C.3 gives that
$\mathcal{P}^{\prime}(E)=\Gamma(E,\mathbb{R}^{n})=\sum_{\tilde{z}\in S}\Gamma((E,\mathbb{R}^{n})_{\tilde{z}},\tilde{z})=\sum_{\tilde{z}\in S}\Gamma(E,\tilde{z})=\sum_{\tilde{z}\in S}\Gamma_{\tilde{z}}(E)\mathcal{Q}(\tilde{z}).$
@@
$\mathcal{P}^{\prime}(\cdot)=\sum_{\tilde{z}\in S}\Gamma_{\tilde{z}}(\cdot)\mathcal{Q}(\tilde{z}).$
@@
where the last line holds as $\mathcal{Q}(\tilde{z})$ is a constant. Hence
@@
$\mathcal{P}(E)=$ $\,\mathcal{P}(E\cap B_{\theta}(\tilde{z}))+\mathcal{P}(E\cap B_{\theta}^{c}(\tilde{z}))$
$=$ $\,\mathcal{P}(E\cap B_{\theta}(\tilde{z}))+\mathcal{P}(E\cap B_{\theta}^{c}(\tilde{z}))+(1-2\delta^{\prime})\mathcal{Q}(\tilde{z})\Gamma_{\tilde{z}}(E)-(1-2\delta^{\prime})\mathcal{Q}(z^{*})\Gamma_{\tilde{z}}(E)$
$=$ $\,\mathcal{P}(E\cap B_{\theta}(\tilde{z}))-(1-2\delta^{\prime})\mathcal{Q}(\tilde{z})\Gamma_{\tilde{z}}(E\cap B_{\theta}(\tilde{z}))$
$+\mathcal{P}(E\cap B_{\theta}^{c}(\tilde{z}))-(1-2\delta^{\prime})\mathcal{Q}(\tilde{z})\Gamma_{\tilde{z}}(E\cap B_{\theta}^{c}(\tilde{z}))$
@@
$c_{\tilde{z},\mathsf{mid}}=\mathcal{P}(B_{\theta}(\tilde{z}))-(1-2\delta^{\prime})\mathcal{Q}(\tilde{z})\Gamma_{\tilde{z}}(B_{\theta}(\tilde{z})),\quad c_{\tilde{z},\mathsf{ext}}=\mathcal{P}(B_{\theta}^{c}(\tilde{z}))-(1-2\delta^{\prime})\mathcal{Q}(\tilde{z})\Gamma_{\tilde{z}}(B_{\theta}^{c}(\tilde{z})).$
@@
$\mathcal{P}(E)\geq(1-2\delta^{\prime})\mathcal{P}^{\prime}(E)=(1-2\delta^{\prime})\sum_{\tilde{z}\in S}\Gamma_{\tilde{z}}(E)\mathcal{Q}(\tilde{z})\geq(1-2\delta^{\prime})\Gamma_{\tilde{z}}(E)\mathcal{Q}(\tilde{z}).$
@@
$\mathrm{supp}(\mathcal{P}_{\tilde{z},\mathsf{mid}})\subseteq B_{\theta}(\tilde{z})\quad\text{and}\quad\mathrm{supp}(\mathcal{P}_{\tilde{z},\mathsf{ext}})\subseteq B^{c}_{\theta}(\tilde{z}),$
@@
$\mathrm{supp}(\mathcal{P}_{\tilde{z},\mathsf{int}})\subseteq B_{\eta^{\prime}}(\tilde{z})\subseteq B_{\theta}(\tilde{z}).$
@@
$=\frac{\mathbb{P}(z^{*}\sim\Gamma_{\tilde{z}})}{(1-2\delta^{\prime})}\frac{1}{\mathbb{P}(z^{*}\sim\Gamma_{\tilde{z}})}\int\int 1_{E}1_{z^{*}\sim\Gamma_{\tilde{z}}}\,\mathrm{d}\mathcal{P}(z^{*})\,\mathrm{d}\mathcal{P}(\cdot|A,u)(\hat{z}).$
Recall that $z^{*}\sim\Gamma_{\tilde{z}}$ is supported in $B_{\eta^{\prime}}(\tilde{z})$. Therefore, for the event $E$ to occur, i.e., ${\|z^{*}-\hat{z}\|}>(c+1)(\eta+\sigma)$, it must be that $\hat{z}\in B_{\theta}^{c}(\tilde{z})$, which means that $\hat{z}\sim\mathcal{P}_{\tilde{z},\mathsf{ext}}(\cdot|A,u)$. Hence
$\mathcal{Q}(\tilde{z})\mathbb{E}_{z^{*}\sim\Gamma_{\tilde{z}},\hat{z}\sim\mathcal{P}(\cdot|A,u)}[1_{E}]$ $\leq\frac{1}{1-2\delta^{\prime}}\int\int 1_{\hat{z}\sim\mathcal{P}_{\tilde{z},\mathsf{ext}}(\cdot|A,u)}1_{z^{*}\sim\Gamma_{\tilde{z}}}\,\mathrm{d}\mathcal{P}(z^{*})\,\mathrm{d}\mathcal{P}(\cdot|A,u)(\hat{z})$
@@
Now let $\mathcal{H}_{\tilde{z},A}$ be the distribution of $y=Az+e$ for $z\sim\mathcal{P}$ and $e\sim\mathcal{E}$ independently. Then Lemma C.4 (with $\mathcal{H}=\mathcal{H}_{\tilde{z},A}$, $\mathcal{H}_{1}=\mathcal{H}_{\tilde{z},\mathsf{int},A}$, $\mathcal{H}_{2}=\mathcal{H}_{\tilde{z},\mathsf{mid},A}$, $\mathcal{H}_{3}=\mathcal{H}_{\tilde{z},\mathsf{ext},A}$ and $a_{1}=(1-2\delta^{\prime})\mathcal{Q}(\tilde{z})$, $a_{2}=c_{\tilde{z},\mathsf{mid}}$, $a_{3}=c_{\tilde{z},\mathsf{ext}}$) gives
@@
Now recall that $\mathcal{H}_{\tilde{z},\mathsf{int},A}$ is the pushforward of a measure $\mathcal{P}_{\tilde{z},\mathsf{int}}$ supported in $B_{\eta^{\prime}}(\tilde{z})$, where $\eta^{\prime}=\eta+\varepsilon^{\prime}\leq\eta+\sigma$ and $\mathcal{H}_{\tilde{z},\mathsf{ext},A}$ is the pushforward of a measure $\mathcal{P}_{\tilde{z},\mathsf{ext}}$ supported in $B^{c}_{\theta}(\tilde{z})$, where $\theta=c(\eta+\sigma)\geq\frac{c}{2}(\eta^{\prime}+\sigma)$. Therefore, Lemma C.5 (with $c$ replaced by $c/2$) gives that

### v2

@@ C.2.2 Disjointly-supported measures induce well-separated measurement distributions
The following lemma pertains to the pushforwards of measures supported in $\mathbb{R}^{n}$ via the forward operator $A$ and noise $e$. Specifically, it states that if two distributions $\mathcal{P}_{\mathsf{int}}$ and $\mathcal{P}_{\mathsf{ext}}$ are disjointly supported then their corresponding pushforwards $\mathcal{H}_{\mathsf{int},A}$ and $\mathcal{H}_{\mathsf{ext},A}$ are, on average with respect to $A\sim\mathcal{A}$, well-separated, in the sense of their TV-distance. It is generalization of [49, Lem. 3.2] that allows for arbitrary distributions $\mathcal{A}$ of the forward operators, as opposed to just distributions of Gaussian random matrices.
@@
[…]
@@
[…]
@@ C.2.3 Replacing the real distribution with the approximate distribution
We next establish a result that allows one to upper bound the failure probability based on draws from the real distribution $\mathcal{R}$ with the failure probability based on draws from the approximate distribution $\mathcal{P}$. This lemma is a key technical step that aligns the prior distribution with the posterior. The specific bound is given in terms of the Wasserstein distance between $\mathcal{R}$ and $\mathcal{P}$ and several of the concentration bounds defined in §2. This is a significant generalization of [49, Lem. 3.3] that allows for arbitrary distributions $\mathcal{A}$, $\mathcal{E}$ for the forwards operator and noise.
@@
Let $\varepsilon,\sigma,d,t\geq 0$, $c\geq 1$, $\mathcal{E}$ be a distribution on $\mathbb{R}^{m}$ and $\mathcal{R},\mathcal{P}$ be distributions on $\mathbb{R}^{n}$ such that $W_{\infty}(\mathcal{R},\mathcal{P})\leq\varepsilon$. Let $\Pi$ be an $W_{\infty}$-optimal coupling of $\mathcal{R}$ and $\mathcal{P}$ and define the set $D=\{x^{*}-z^{*}:(x^{*},z^{*})\in\mathrm{supp}(\Pi)\}$. Let
@@
$q=\mathbb{P}_{z^{*}\sim\mathcal{P},A\sim\mathcal{A},e\sim\mathcal{E},\hat{z}\sim\mathcal{P}(\cdot|Az^{*}+e,A)}[{\|z^{*}-\hat{z}\|}\geq d].$
@@
[…]
@@
$\times\int 1_{B_{2,\hat{x}}}(z^{*})\,\mathrm{d}\mathcal{P}(\cdot|Az^{*}+e^{\prime},A)(\hat{x})\,\mathrm{d}e^{\prime}\,\mathrm{d}\mathcal{A}(A)\,\mathrm{d}\Pi(x^{*},z^{*}).$
@@ C.2.4 Decomposing distributions
The following lemma is in large part similar to [49, Lem. A.1]. However, we streamline and rewrite its proof for clarity and completeness, fix a number of small issues and make an addition to the statement (see item (v) below) that is important for proving our main result.
@@
then there exist distributions $\mathcal{R}^{\prime},\mathcal{R}^{\prime\prime},\mathcal{P}^{\prime},\mathcal{P}^{\prime\prime}$, a constant $0<\delta^{\prime}\leq\delta$ and a discrete distribution $\mathcal{Q}$ with $\mathrm{supp}(\mathcal{Q})=S$ satisfying
@@
This lemma states that two distributions that are close in Wasserstein $p$-distance, and for which at least one has small approximate covering number (C.15), can be decomposed into mixtures (iii) of distributions, where the following holds. One of the distributions, say $\mathcal{P}^{\prime}$, is close (i) in Wasserstein-$\infty$ distance to a discrete distribution $\mathcal{Q}$ with the cardinality of its support (iv) bounded by the approximate covering number. The other $\mathcal{R}^{\prime}$ is close in Wasserstein-$\infty$ distance to $\mathcal{P}^{\prime}$. Moreover, both mixtures (iii) are dominated by these distributions: the ‘remainder’ terms $\mathcal{P}^{\prime\prime}$ and $\mathcal{R}^{\prime\prime}$ are associated with a small constant $\delta^{\prime}\leq\delta$, meaning they are sampled with probability $\leq\delta$ when drawing from either $\mathcal{P}$ or $\mathcal{R}$. Note that if $p<\infty$ then the Wasserstein-$\infty$ distance between $\mathcal{R}^{\prime}$ and $\mathcal{P}^{\prime}$ may get larger as $\delta$ shrinks, i.e., as the remainder gets smaller. However, this does not occur when $p=\infty$, as (ii) is independent of $\delta$ in this case.
@@
[…]
@@
[…]
@@
[…]
@@
We have measures $\mathcal{P}$, $\mathcal{Q}^{\prime}$ and $\mathcal{R}$ and couplings $\Pi$ of $\mathcal{P},\mathcal{Q}^{\prime}$ and $\Gamma$ of $\mathcal{P}$, $\mathcal{R}$. We will in a sense, couple $\Pi,\Gamma$. Since $(\mathbb{R}^{n})^{3}$ is a Polish space, by [10, Lem. 8.4], there exists a coupling $\Omega$ with
@@
$E_{1}$ $:=\{(x_{1},x_{2},x_{3}):{\|x_{1}-x_{2}\|}>\eta,z\in\mathbb{R}^{n}\},$
$E_{2}$ $:=\{(x_{1},x_{2},x_{3}):{\|x_{1}-x_{3}\|}>\rho/\delta^{1/p},y\in\mathbb{R}^{n}\},$
@@
$\mathcal{R}^{\prime\prime}(A)$ $=\Omega(\mathbb{R}^{n},\mathbb{R}^{n},A|E^{c}),$
@@
We now prove Theorem 3.1. This follows a similar approach to that of [49, Thm. 3.4], but with a series of significant modifications to account for the substantially more general setup considered in this work. We also streamline the proof and clarify a number of key steps.
@@
[…]
@@
[…]
@@
[…]
@@
$\mathcal{P}(E)=$ $)\displaystyle\,\mathcal{P}(E\cap B_{\theta}(\tilde{z}))+\mathcal{P}(E\cap B_{\theta}^{c}(\tilde{z}$))
$=$ $\,\mathcal{P}(E\cap B_{\theta}(\tilde{z}))+\mathcal{P}(E\cap B_{\theta}^{c}(\tilde{z}))+(1-2\delta^{\prime})\mathcal{Q}(\tilde{z})\Gamma_{\tilde{z}}(E)-(1-2\delta^{\prime})\mathcal{Q}(\tilde{z})\Gamma_{\tilde{z}}(E)$
$=$ $)\displaystyle\,\mathcal{P}(E\cap B_{\theta}(\tilde{z}))-(1-2\delta^{\prime})\mathcal{Q}(\tilde{z})\Gamma_{\tilde{z}}(E\cap B_{\theta}(\tilde{z}$))
$)\displaystyle+\mathcal{P}(E\cap B_{\theta}^{c}(\tilde{z}))-(1-2\delta^{\prime})\mathcal{Q}(\tilde{z})\Gamma_{\tilde{z}}(E\cap B_{\theta}^{c}(\tilde{z}$))
@@
$).\displaystyle c_{\tilde{z},\mathsf{mid}}=\mathcal{P}(B_{\theta}(\tilde{z}))-(1-2\delta^{\prime})\mathcal{Q}(\tilde{z})\Gamma_{\tilde{z}}(B_{\theta}(\tilde{z})),\quad c_{\tilde{z},\mathsf{ext}}=\mathcal{P}(B_{\theta}^{c}(\tilde{z}))-(1-2\delta^{\prime})\mathcal{Q}(\tilde{z})\Gamma_{\tilde{z}}(B_{\theta}^{c}(\tilde{z}$)).
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]